OpenStudy (anonymous):

Help me out From 5 women nd 7 men, a committee of 2 women and 3 men is to be formed (a) same number of persons but two men refused to work together (b) rashed nd his girlfrnd always wants to work together

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi

OpenStudy (misty1212):

\[\binom{5}{2}=\frac{5\times 4}{2}=10\]choices for the two women

OpenStudy (anonymous):

Can u explain @misty1212

OpenStudy (misty1212):

i am still thinking about the second part, 3 out of 7 men but two refuse to serve together

OpenStudy (anonymous):

See i have solve it but donnow if its correct or not

OpenStudy (misty1212):

if there were no restrictions on the men it would be \[\binom{5}{2}\times \binom{7}{3}\] but we need to figure out how many ways the two men could be on the same committee so we can subtract that off

OpenStudy (anonymous):

5C2[(5C2×2C1)+(5C3×2C0)]= 300

OpenStudy (mathmate):

@misty1212 The two men who refuse to work together can be considered a single person, one or the other, or none at all.

OpenStudy (misty1212):

so \(\binom{6}{3}\)?

OpenStudy (mathmate):

yes!

OpenStudy (anonymous):

@mathmate check my answer and methid

OpenStudy (anonymous):

@misty1212 check my method

OpenStudy (anonymous):

@mathmate check my answer and methid

OpenStudy (anonymous):

5C2[(5C2×2C1)+(5C3×2C0)]= 300

OpenStudy (anonymous):

See i have solve it but donnow if its correct or not

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Can u explain @misty1212

OpenStudy (mathmate):

@misty1212 oops, on second thought, (6,3) won't work, because we would be missing something. It would be more like (5,3)+2(5,2) instead of (7,3).

OpenStudy (mathmate):

@statisz Are there three parts to the question? the question itself, then (a), and (b)?

OpenStudy (anonymous):

I answered part a

OpenStudy (mathmate):

@statisz The question is not very clear to me. Are there three parts to the question? the question itself, then (a), and (b)? and Is the condition in (a) (where the two men refuse to work together) included in (b)?

OpenStudy (anonymous):

No a is not included in b

OpenStudy (mathmate):

@statisz The question without constraints (a) and (b) has been answered by @misty1212, i.e. (5,2)*(7,3), namely the product of number of ways of choosing the women (5 choose 2) and number of ways of choosing the men (7 choose 3). ok so far?

OpenStudy (anonymous):

Yup okay

OpenStudy (anonymous):

And dn if two mens refuse to be togther dn wat to do

OpenStudy (anonymous):

@mathmate help.me plz Or tag someone who knows this

OpenStudy (mathmate):

For part (b) there are two cases, either 1. Rached and his gf are in: then there are 1 woman and 2 other men left in the choices, namely (4,1)*(6,2), or 2. Rached and his gf are out: then there are 4 women and 6 men to choose from, making (4,2)*(6,3) Add the two cases together to give part (b). Part (a) will follow.

OpenStudy (mathmate):

For case (a): Start with (5,3), i.e. not choosing any of the trouble makers, then \(add\) (2,1)*(5,2) by choosing one of the problematic men plus 2 other men. The total is then (5,3) + 2(5,2)=30. Another way to work is to include one of the problematic men, then there are (6,3)=20 ways. Including the 7th man would add C(1,1)*(5,2)=10 for a total of 30 again.

OpenStudy (anonymous):

Tnx @mathmate